OpenStudy (anonymous):

Find dy/dx by implicit differentiation. 3x^2 + 5xy - y^2 = 8

OpenStudy (amistre64):

implicit differentation: (d/dx)(3x^2) +(d/dx)(5xy) -(d/dx)(y^2) = (d/dx)(8) dy/dx = y' 6x + (5xy' + 5y) -2yy' =0 y'(5x -2y) +6x +5y=0 y'(5x -2y) = -6x -5y y' = (-6x -5y)/(5x -2y) That should be it :)

OpenStudy (anonymous):

thank you!